# Optimizing Quantum Transformation Matrices: A Block Decomposition Approach for Efficient Gate Reduction

### v1

@@ IV.2 Circuit Formulation
By optimizing $\mathbf{Y}_{0}$, the resultant matrix $\mathbf{Y}$, constructed with $10$ gates, predicts the state in Eq. (21) as:
$\begin{split}\mathbf{Y}|W\rangle=-0.142|000\rangle+0.056|001\rangle-0.088|010\rangle+0.180|011\rangle\\ +0.562|100\rangle+0.072|101\rangle+0.388|110\rangle+0.007|111\rangle\end{split}$ (23)
@@
The comparison between the original transformation $\mathbf{U}|W\rangle$ and the approximations, $\mathbf{Y}_{0}|W\rangle$ and $\mathbf{Y}|W\rangle$ showcases the impact of gate optimization on accuracy. Extracting 10 gates randomly from $\mathbf{U}$ to form $\mathbf{Y}_{0}$ yields a notably poor estimate, deviating significantly from the true transformation $\mathbf{U}|W\rangle$. However, through optimization using the proposed algorithm, the refined transformation matrix $\mathbf{Y}$ constructed with the same number of gates remarkably aligns closer with $\mathbf{U}|W\rangle$. This enhancement underscores the merit gains from optimization in approximating complex unitary transformations.
V Conclusion
@@
In conclusion, our paper presented an alternative method for approximating complex unitary transformations with a focus on limiting the number of gates. By considering the block decomposition technique, we successfully sparsed the transformation matrix with an adjustable gate count choice and achieved good approximation accuracy. Through simulations involving a 3-qubit transformation, we demonstrated the effectiveness of our approach and showcased the trade-offs between accuracy and computational complexity. Our study not only highlighted the manageable computational complexity but also underscored the limitations as the quantum system scales up in qubit numbers.
@@ Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant No. 12474489), Shenzhen Fundamental Research Program (Grant No. JCYJ20240813153139050), the Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2203001, GDZX2403001), and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302300).
@@ Appendix A Quantum Gate Positioning and Transformation Matrix Interpretation
$\begin{array}[]{c|cccccccc}&|000\rangle&|001\rangle&|010\rangle&|011\rangle&|100\rangle&|101\rangle&|110\rangle&|111\rangle\\ \hline\cr|000\rangle&u_{1,1}&u_{1,2}&u_{1,3}&u_{1,4}&u_{1,5}&u_{1,6}&u_{1,7}&u_{1,8}\\ |001\rangle&u_{2,1}&u_{2,2}&u_{2,3}&u_{2,4}&u_{2,5}&u_{2,6}&u_{2,7}&u_{2,8}\\ |010\rangle&u_{3,1}&u_{3,2}&u_{3,3}&u_{3,4}&u_{3,5}&u_{3,6}&u_{3,7}&u_{3,8}\\ |011\rangle&u_{4,1}&u_{4,2}&u_{4,3}&u_{4,4}&u_{4,5}&u_{4,6}&u_{4,7}&u_{4,8}\\ |100\rangle&u_{5,1}&u_{5,2}&u_{5,3}&u_{5,4}&u_{5,5}&u_{5,6}&u_{5,7}&u_{5,8}\\ |101\rangle&u_{6,1}&u_{6,2}&u_{6,3}&u_{6,4}&u_{6,5}&u_{6,6}&u_{6,7}&u_{6,8}\\ |110\rangle&u_{7,1}&u_{7,2}&u_{7,3}&u_{7,4}&u_{7,5}&u_{7,6}&u_{7,7}&u_{7,8}\\ |111\rangle&u_{8,1}&u_{8,2}&u_{8,3}&u_{8,4}&u_{8,5}&u_{8,6}&u_{8,7}&u_{8,8}\\ \end{array}$ (24)

### v2

@@ IV.2 Circuit Formulation
Optimizing $\mathbf{Y}_{0}$, the resultant matrix $\mathbf{Y}$, constructed with $10$ gates, predicts the state in Eq. (27) as:
$\begin{split}\mathbf{Y}|W\rangle=&\left(0\right)|000\rangle+\left(-0.036-0.089i\right)|001\rangle\\ &+\left(0.015-0.109i\right)|010\rangle+\left(-0.058-0.018i\right)|011\rangle\\ &+\left(-0.178+0.125i\right)|100\rangle+\left(0.349+0.222i\right)|101\rangle\\ &+\left(-0.125-0.099i\right)|110\rangle+\left(0.191+0.419i\right)|111\rangle\end{split}$ (29)
To demonstrate the effectiveness of our approach, we evaluate the fidelity of both the resulting matrices $\mathbf{Y}_{0}$ and $\mathbf{Y}$ with respect to the target matrix $\mathbf{U}$. The fidelity can be calculated by introducing the density matrices $|\psi\rangle=\mathbf{Y}|W\rangle$ and $|\phi\rangle=\mathbf{U}|W\rangle$, and then applying the fidelity formula:
$F(|\psi\rangle,|\phi\rangle)=|\langle\psi|\phi\rangle|^{2}$ (30)
Since our method can potentially be applied to any quantum operator, such as sparse unitary transformation matrices produced by the method in Man and Wang (2024), applying these operators may lead to unnormalized states due to approximations. The comparison between the original transformation $\mathbf{U}|W\rangle$ and the approximations $\mathbf{Y}_{0}|W\rangle$ and $\mathbf{Y}_{0}|W\rangle$ showcases the impact of gate optimization on accuracy. To ensure meaningful comparisons, we first normalize the states $\mathbf{U}|W\rangle$ (27), $\mathbf{Y}_{0}|W\rangle$ (28), and $\mathbf{Y}_{0}|W\rangle$ (29) as follows:
$|\psi_{\text{normalized}}\rangle=\frac{1}{\sqrt{\sum_{n}|c_{n}|^{2}}}\sum_{n}c_{n}|n\rangle.$ (31)
By applying (31) to $\mathbf{U}|W\rangle$ (27), $\mathbf{Y}_{0}|W\rangle$ (28), and $\mathbf{Y}_{0}|W\rangle$ (29), the normalized states are given as:
$\begin{split}U|W\rangle_{\text{normalized}}=&\frac{1}{0.9702}\big[(-0.019+0.0191i)|000\rangle+(-0.171-0.254i)|001\rangle\\ &+(-0.099-0.602i)|010\rangle+(-0.058-0.018i)|011\rangle\\ &+(0.293+0.163i)|100\rangle+(0.253+0.139i)|101\rangle\\ &+(-0.232-0.031i)|110\rangle+(0.161+0.441i)|111\rangle\big],\end{split}$ (32)
$\begin{split}Y_{0}|W\rangle_{\text{normalized}}=&\frac{1}{0.9995}\big[(-0.012+0.026i)|000\rangle+(-0.020+0.090i)|001\rangle\\ &+(-0.528+0.005i)|010\rangle+(0.488-0.188i)|011\rangle\\ &+(-0.043-0.100i)|100\rangle+(-0.136+0.345i)|101\rangle\\ &+(0.454+0.153i)|110\rangle+(-0.196+0.142i)|111\rangle\big],\end{split}$ (33)
$\begin{split}Y|W\rangle_{\text{normalized}}=&\frac{1}{0.6935}\big[(0)|000\rangle+(-0.036-0.089i)|001\rangle\\ &+(0.015-0.109i)|010\rangle+(-0.058-0.018i)|011\rangle\\ &+(-0.178+0.125i)|100\rangle+(0.349+0.222i)|101\rangle\\ &+(-0.125-0.099i)|110\rangle+(0.191+0.419i)|111\rangle\big].\end{split}$ (34)
The fidelity between the normalized states is calculated as follows:
$F(Y_{0}|W\rangle,U|W\rangle)=|\langle Y_{0}|W_{\text{normalized}}|U|W_{\text{normalized}}\rangle|^{2}\approx 0.853.$ (35)
$F(Y|W\rangle,U|W\rangle)=|\langle Y|W_{\text{normalized}}|U|W_{\text{normalized}}\rangle|^{2}\approx 0.921.$ (36)
By using Eq. (30), the fidelity of $\mathbf{Y}_{0}$ and $\mathbf{U}$ is $0.853$ from  (35), whereas the fidelity between $\mathbf{Y}$ and $\mathbf{U}$ is $0.921$ from  (36). Normalization ensures that the states represent valid quantum states, and the fidelity calculation reflects the true overlap between the states. Since fidelity measures the closeness of two quantum states, higher fidelity indicates that the two states are more similar, this showed that $\mathbf{Y}$ yield a better result.
@@
This comparison demonstrates that our optimization process significantly improves the approximation of the target transformation. Extracting 10 gates randomly from $\mathbf{U}$ to form $\mathbf{Y}_{0}$ yields a notably poor estimate, deviating significantly from the true transformation $\mathbf{U}|W\rangle$. However, through optimization using the proposed algorithm, the refined transformation matrix $\mathbf{Y}$ constructed with the same number of gates is remarkably aligned closer to $\mathbf{U}|W\rangle$. This enhancement underscores the merit gains from optimization in the approximation of complex unitary transformations.
V Discussion
Our work builds upon and extends several existing methods in quantum circuit optimization and gate reduction. Traditional approaches to quantum circuit decomposition, such as those based on the Cosine-Sine Decomposition method Mottonen et al. (2004); Di and Wei (2013), focus on breaking down unitary matrices into sequences of elementary gates. Although these methods are effective in simplifying quantum circuits, they often result in a large number of gates, especially for high-dimensional transformations. In contrast, our approach introduces flexibility by allowing users to specify the desired number of gates, thereby optimizing gate usage while maintaining computational accuracy. This is particularly advantageous in scenarios where the gate count is a critical resource constraint.
In the field of optimization, recent work by Yuan et al. (2020) introduced the block decomposition algorithm, which has shown remarkable success in sparse optimization problems. Our method adapts this approach specifically for quantum gate reduction by incorporating unitary constraints and gate structure constraints. This adaptation allows us to achieve a balance between gate count reduction and transformation accuracy, which is not explicitly addressed in the original block decomposition framework. Furthermore, our work differs from gate reduction strategies that rely on automated optimization techniques, such as those discussed in Nam et al. (2018) and Abdessaied et al. (2014). Although these methods aim to reduce gate counts through iterative refinement and library-based optimization, they do not provide explicit control over the number of gates used in the final circuit. Our algorithm, on the other hand, allows users to directly specify the maximum number of gates, offering a more flexible and resource-efficient approach to quantum circuit design.
The significance of our work lies in its ability to bridge the gap between sparse optimization techniques and quantum circuit design. By integrating the block decomposition method with unitary constraints, we provide an alternative approach to approximating quantum transformations with a limited number of gates. This integration introduces flexibility in resource allocation, enabling users to tailor the gate count to specific hardware limitations or application requirements.
This flexibility represents a significant improvement over traditional decomposition methods, which typically produce circuits with a fixed number of gates based on the dimensionality of the transformation matrix. In summary, our work contributes to the growing body of research on quantum circuit optimization by introducing a new method that combines the strengths of block decomposition with the specific constraints of quantum gate design. This approach not only enhances the efficiency of quantum computations but also provides a more adaptable framework for designing quantum circuits in resource-constrained environments.
VI Conclusion
@@
In conclusion, our paper presented an alternative method for approximating complex unitary transformations with a focus on limiting the number of gates. Considering the block decomposition technique, we successfully sparsed the transformation matrix with an adjustable gate count choice and achieved good approximation accuracy. Through simulations involving a 3-qubit transformation, we demonstrated the effectiveness of our approach and showcased the trade-offs between accuracy and computational complexity. Our study not only highlighted the manageable computational complexity but also underscored the limitations as the quantum system scales up in qubit numbers.
@@ Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant No. 11874312, 12474489), Shenzhen Fundamental Research Program (Grant No. JCYJ20240813153139050), the Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2203001, GDZX2403001), and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302300).
The data that support the findings of this article are openly available dataavail .
@@ Appendix A Quantum Gate Positioning and Transformation Matrix Interpretation
[…]